OpenStudy (anonymous):

if you travel to work at 30mph and go home by 40 mph, what is the average rate for the entire trip?

OpenStudy (anonymous):

I am totally lost, because there is no distance or time given

OpenStudy (anonymous):

are you there? I appreciate the help

OpenStudy (anonymous):

That's all the problem gives you?

OpenStudy (anonymous):

yes, it is why I am so frustated

OpenStudy (regan):

(30mph*t1+40mph*t2)/(t1+t2) which reads total distance over total time thats all i can think of

OpenStudy (anonymous):

yeah, regan's answer looks right

OpenStudy (anonymous):

if you aren't given time or distance

OpenStudy (regan):

and 30mph*t1=40mph*t2 this may also help

OpenStudy (regan):

30mph*t1*2/t1+30mph*t1/40mph

OpenStudy (regan):

t1(30*2) numerator t1(1+3/4) denominator

OpenStudy (regan):

34.29 mph : )

OpenStudy (anonymous):

how'd you get that lol :)

OpenStudy (regan):

ok using 30mph*t1=40mph*t2 i said 30mph*t1+30mph*t1=30mph*t1+40mph*t2 in the denominator and t1+t2=t1+(30/40)*t1 in the numerator so we have (t1(30*2))/(t1(1+3/4)) numerator/denominator however they both have been written in terms of t1 therefore can cancel to get 30*2/1.75 which equals 34.29

OpenStudy (regan):

does that make sense?

OpenStudy (anonymous):

yeah it does , thank you so much